OpenStudy (anonymous):

Change the given equation into slope-intercept form. 3x + 4y = 12

OpenStudy (anonymous):

Y=3/4X+3....BANG BANG [:

OpenStudy (anonymous):

sosa :))

OpenStudy (anonymous):

Standard form. Ax + By = C Slope-Intercept form. y = mx + b You need to solve for y. \[ 3x + 4y = 12 \]Subtract 3x from both sides.\[4y = -3y + 12\]Divide both sides by 4.\[y = \frac{ -3 }{4 } x + 3\]

OpenStudy (anonymous):

told u ^^ ;D

OpenStudy (anonymous):

opps forgot the (-) buh yae

OpenStudy (anonymous):

:)) thank u girl:)) sosa :)) banh bang

OpenStudy (anonymous):

wepa ;D